At 6 A.M. a snowplow, traveling at a constant speed, begins to clear a highway leading out of town. At 8 A.M. an automobile begins traveling the highway at a speed of 25 mi/hr and reaches the plow 30 minutes later. Find the speed of the snowplow.
Answer by richwmiller(17219)

The snowplow had a two head start and went 2x
The car gained on the plow at (25-x) every hour
x and 25
2x/(25-x)=1/2
4x=25-x
5x=25
x=5 mph
let's check
5 mph for the plow
In two hours the plow went 10 miles
The car gains at 25-5=20 mph
10/20=1/2
0k